# Hi-SLAM: Scaling-up Semantics in SLAM with a Hierarchically Categorical Gaussian Splatting

### v2

@@ Abstract
We propose Hi-SLAM, a semantic 3D Gaussian Splatting SLAM method featuring a novel hierarchical categorical representation, which enables accurate global 3D semantic mapping, scaling-up capability, and explicit semantic label prediction in the 3D world. The parameter usage in semantic SLAM systems increases significantly with the growing complexity of the environment, making it particularly challenging and costly for scene understanding. To address this problem, we introduce a novel hierarchical representation that encodes semantic information in a compact form into 3D Gaussian Splatting, leveraging the capabilities of large language models (LLMs). We further introduce a novel semantic loss designed to optimize hierarchical semantic information through both inter-level and cross-level optimization. Furthermore, we enhance the whole SLAM system, resulting in improved tracking and mapping performance. Our Hi-SLAM outperforms existing dense SLAM methods in both mapping and tracking accuracy, while achieving a 2x operation speed-up. Additionally, it exhibits competitive performance in rendering semantic segmentation in small synthetic scenes, with significantly reduced storage and training time requirements. Rendering FPS impressively reaches 2,000 with semantic information and 3,000 without it. Most notably, it showcases the capability of handling the complex real-world scene with more than 500 semantic classes, highlighting its valuable scaling-up capability.
@@ I INTRODUCTION
Figure 1: (a). The global 3D Gaussian map generated by Hi-SLAM with learned semantic labels is shown on the left. The hierarchical structure of the semantic information is organized on the right, considering both semantic and geometric attributes (blue box). The proposed hierarchical categorical representation compresses semantic data, reducing both memory usage and training time of the semantic SLAM. (b). The rendered semantic map at different levels shows a coarse-to-fine understanding, beneficial for real-world scenarios with shifting perspectives from distant to close.
@@
Thus, it is highly desirable to extend the original 3D Gaussian Splatting with semantic capabilities while preserving its advantageous probabilistic representation. A straightforward approach would be to augment 3D points with a discrete semantic label and parameterize its distribution with a categorical discrete distribution, i.e., a flat Softmax embedding representation. However, 3D Gaussian Splatting is already a storage-intensive representation [14, 15], requiring a large number of 3D primitives with multiple parameters to achieve realistic rendering. Adding semantic distribution parameters would result in significantly increased storage demands and processing time, growing linearly with the number of semantic classes. This makes it particularly impractical for complex scene understanding. Recent works formulate semantic classes using non-distributional approaches to handle this complexity. The work [16] directly learns a 3-channel RGB visualization for semantic maps instead of the true semantic information understanding. Another work [17] uses a flat semantic representation with supervision from pre-trained foundation models, yet it can only produces a 3D semantic feature map, rather than an explicit 3D semantic label map.
@@
Building on this concept, we propose Hi-SLAM, a Semantic Gaussian Splatting SLAM leveraging the hierarchical categorical representation for semantic information. Specifically, taking both semantic and geometric attributes into consideration, a well-designed tree is established with the help of Large Language Models (LLMs), which significantly reduces memory usage and training time, effectively compressing data while preserving its physical meaning. Additionally, we introduce a hierarchical loss for the proposed representation, incorporating both inter-level and cross-level optimizations. This strategy facilitates a coarse-to-fine understanding of scenes, which aligns well with real-world applications, particularly those involving observations from distant to nearby views. Furthermore, we enhance and refine the Gaussian SLAM to improve both performance and running speed.
@@
3) We update several modules to enhance the SLAM system by incorporating semantic information and fully leveraging the rendering capabilities of Gaussian Splatting.
@@
Figure 2: Left: Overview of the Hi-SLAM pipeline. The global 3D Gaussian map is initialized with the first image. The system then alternates between Tracking and Mapping steps as new frames are processed (see Section III-C). Top Right: Hierarchical representation of semantic information. The Tree Generation process uses a Loop-based critic operation, including a LLM and a Validator, to create a tree coding from leaf-to-root. This tree is used to establish hierarchical coding for each Gaussian primitive (see Section III-A). Additionally, a novel loss combining Inter-level Loss $L_{\text{Inter}}$ and Cross-level Loss $L_{\text{Cross}}$ is proposed for hierarchical semantic optimization (see Section III-B). Bottom Right: An example of hierarchical semantic rendering.
@@ III-C Gaussian Splatting Semantic Mapping and Tracking
The pipeline of our Hi-SLAM is illustrated in Fig. 2. We will detail the submodules in this subsection.
@@ IV-B SLAM Performance
Running time. Running times for all methods are shown in Tab. IV. Compared to state-of-the-art dense visual SLAM approaches, our method (Ours*) achieves up to 2.4× faster tracking and 2.2× faster mapping than the SOTA performance [9]. When incorporating semantic information, our method remains efficient, leveraging hierarchical semantic coding to achieve nearly 3× faster tracking and 1.2× faster mapping compared with the semantic SLAM with flat semantic coding. Notably, our Hi-SLAM achieves a rendering speed of 2000 FPS. For Hi-SLAM without semantic information, the rendering speed increases to 3000 FPS.
@@
Hi-SLAM (Ours*) 18.71 22.93 0.75 1.38
Hi-SLAM (Ours) 46.90 148.66 1.88 8.92
Hi-SLAM (Ours) 61.23 170.30 2.45 10.22
Hi-SLAM (Ours**) 168.94 204.25 6.75 12.26
@@ IV-C Hierarchical semantic understanding
Hi-SLAM (Ours**) 90.35 91.21 90.62 89.11 90.45
Hi-SLAM (Ours) 76.44 76.62 78.31 80.39 70.43
Param (MB) Hi-SLAM (Ours**) 2662.25 2355 3072 2560 2662
Hi-SLAM (Ours) 910.50 793 1126 843 880
@@
Figure 4: Visualization of the established semantic 3D map across multiple levels, demonstrating a coarse-to-fine semantic understanding of the complex scene. The bottom of the figure displays localization, mapping, and rendering performance, providing a comprehensive overview of Hi-SLAM’s effectiveness.
@@ V CONCLUSIONS
We present Hi-SLAM, a novel semantic 3D Gaussian Splatting SLAM method with a hierarchical categorical representation, which can generate global 3D semantic map with scaling-up capability and explicit semantic semantic label prediction. Specifically, we propose a novel hierarchical representation to encode semantic message in a compact form, further forming it into 3D Gaussian Splatting, with the assistance of LLMs. Additionally, we present a novel semantic loss for the hierarchical semantic information optimization, including both inter-level and cross-level optimization. Furthermore, we refine the whole SLAM system. Experiments demonstrate that Hi-SLAM surpasses existing dense SLAM methods in both tracking and mapping performance with faster speed, and it can significantly reduces storage and training requirements. The method exhibits exceptional rendering performance, achieving up to 2,000/3,000 FPS with/without semantic information. Most notably, Hi-SLAM showcases its ability to manage complex real-world scenes containing over 500 semantic classes, underlining its robust scalability and effectiveness in large-scale applications.

### v3

# Hier-SLAM: Scaling-up Semantics in SLAM with a Hierarchically Categorical Gaussian Splatting

@@ Abstract
We propose Hier-SLAM, a semantic 3D Gaussian Splatting SLAM method featuring a novel hierarchical categorical representation, which enables accurate global 3D semantic mapping, scaling-up capability, and explicit semantic label prediction in the 3D world. The parameter usage in semantic SLAM systems increases significantly with the growing complexity of the environment, making it particularly challenging and costly for scene understanding. To address this problem, we introduce a novel hierarchical representation that encodes semantic information in a compact form into 3D Gaussian Splatting, leveraging the capabilities of large language models (LLMs). We further introduce a novel semantic loss designed to optimize hierarchical semantic information through both inter-level and cross-level optimization. Furthermore, we enhance the whole SLAM system, resulting in improved tracking and mapping performance. Our Hier-SLAM outperforms existing dense SLAM methods in both mapping and tracking accuracy, while achieving a 2x operation speed-up. Additionally, it exhibits competitive performance in rendering semantic segmentation in small synthetic scenes, with significantly reduced storage and training time requirements. Rendering FPS impressively reaches 2,000 with semantic information and 3,000 without it. Most notably, it showcases the capability of handling the complex real-world scene with more than 500 semantic classes, highlighting its valuable scaling-up capability.
@@ I INTRODUCTION
Figure 1: (a). The global 3D Gaussian map generated by Hier-SLAM with learned semantic labels is shown on the left. The hierarchical structure of the semantic information is organized on the right, considering both semantic and geometric attributes (blue box). The proposed hierarchical categorical representation compresses semantic data, reducing both memory usage and training time of the semantic SLAM. (b). The rendered semantic map at different levels shows a coarse-to-fine understanding, beneficial for real-world scenarios with shifting perspectives from distant to close.
@@
Thus, it is highly desirable to extend the original 3D Gaussian Splatting with semantic capabilities while preserving its advantageous probabilistic representation. A straightforward approach would be to augment 3D points with a discrete semantic label and parameterize its distribution with a categorical discrete distribution, i.e., a flat Softmax embedding representation. However, 3D Gaussian Splatting is already a storage-intensive representation [14, 15], requiring a large number of 3D primitives with multiple parameters to achieve realistic rendering. Adding semantic distribution parameters would result in significantly increased storage demands and processing time, growing linearly with the number of semantic classes. This makes it particularly impractical for complex scene understanding. Recent works formulate semantic classes using non-distributional approaches to handle this complexity. The work [16] directly learns a 3-channel RGB visualization for semantic maps instead of the semantic label learning. Another work [17] uses a flat semantic representation with supervision from pre-trained foundation models, yet it can only produces a 3D semantic feature map, rather than an explicit 3D semantic label map.
@@
Building on this concept, we propose Hier-SLAM, a Semantic Gaussian Splatting SLAM leveraging the hierarchical categorical representation for semantic information. Specifically, taking both semantic and geometric attributes into consideration, a well-designed tree is established with the help of Large Language Models (LLMs), which significantly reduces memory usage and training time, effectively compressing data while preserving its physical meaning. Additionally, we introduce a hierarchical loss for the proposed representation, incorporating both inter-level and cross-level optimizations. This strategy facilitates a coarse-to-fine understanding of scenes, which aligns well with real-world applications, particularly those involving observations from distant to nearby views. Furthermore, we enhance and refine the Gaussian SLAM to improve both performance and running speed.
@@
3)
@@
Figure 2: Left: Overview of the Hier-SLAM pipeline. The global 3D Gaussian map is initialized with the first image. The system then alternates between Tracking and Mapping steps as new frames are processed (see Section III-C). Top Right: Hierarchical representation of semantic information. The Tree Generation process uses a Loop-based critic operation, including a LLM and a Validator, to create a tree coding from leaf-to-root. This tree is used to establish hierarchical coding for each Gaussian primitive (see Section III-A). Additionally, a novel loss combining Inter-level Loss $L_{\text{Inter}}$ and Cross-level Loss $L_{\text{Cross}}$ is proposed for hierarchical semantic optimization (see Section III-B). Bottom Right: An example of hierarchical semantic rendering.
@@ III-C Gaussian Splatting Semantic Mapping and Tracking
The pipeline of our Hier-SLAM is illustrated in Fig. 2. We will detail the submodules in this subsection.
@@ IV-B SLAM Performance
Running time. Running times for all methods are shown in Tab. IV. Compared to state-of-the-art dense visual SLAM approaches, our method (Ours*) achieves up to 2.4× faster tracking and 2.2× faster mapping than the SOTA performance [9]. When incorporating semantic information, our method remains efficient, leveraging hierarchical semantic coding to achieve nearly 3× faster tracking and 1.2× faster mapping compared with the semantic SLAM with flat semantic coding. Notably, our Hier-SLAM achieves a rendering speed of 2000 FPS. For Hier-SLAM without semantic information, the rendering speed increases to 3000 FPS.
@@
Hier-SLAM (Ours*) 18.71 22.93 0.75 1.38
Hier-SLAM (Ours) 46.90 148.66 1.88 8.92
Hier-SLAM (Ours) 61.23 170.30 2.45 10.22
Hier-SLAM (Ours**) 168.94 204.25 6.75 12.26
@@ IV-C Hierarchical semantic understanding
Hier-SLAM (Ours**) 90.35 91.21 90.62 89.11 90.45
Hier-SLAM (Ours) 76.44 76.62 78.31 80.39 70.43
Param (MB) Hier-SLAM (Ours**) 2662.25 2355 3072 2560 2662
Hier-SLAM (Ours) 910.50 793 1126 843 880
@@
Figure 4: Visualization of the established semantic 3D map across multiple levels, demonstrating a coarse-to-fine semantic understanding of the complex scene. The bottom of the figure displays localization, mapping, and rendering performance, providing a comprehensive overview of Hier-SLAM’s effectiveness.
@@ V CONCLUSIONS
We present Hier-SLAM, a novel semantic 3D Gaussian Splatting SLAM method with a hierarchical categorical representation, which can generate global 3D semantic map with scaling-up capability and explicit semantic semantic label prediction. Specifically, we propose a novel hierarchical representation to encode semantic message in a compact form, further forming it into 3D Gaussian Splatting, with the assistance of LLMs. Additionally, we present a novel semantic loss for the hierarchical semantic information optimization, including both inter-level and cross-level optimization. Furthermore, we refine the whole SLAM system. Experiments demonstrate that Hier-SLAM surpasses existing dense SLAM methods in both tracking and mapping performance with faster speed, and it can significantly reduces storage and training requirements. The method exhibits exceptional rendering performance, achieving up to 2,000/3,000 FPS with/without semantic information. Most notably, Hi-SLAM showcases its ability to manage complex real-world scenes containing over 500 semantic classes, underlining its robust scalability and effectiveness in large-scale applications.